OpenStudy (anonymous):

Please help? Write an equation of the line, in point-slope form, that passes through the two given points. points: (–2, 15), (9, –18)

OpenStudy (anonymous):

can you calculate the slope of the line passing through those two points? use: \(\Large slope=m=\frac{y_2-y_1}{x_2-x_1} \)

OpenStudy (anonymous):

do you know how to use the slope formula?

OpenStudy (anonymous):

would the answer be y-15=-3(x+2)

OpenStudy (anonymous):

yep...:)

OpenStudy (anonymous):

of course, the answer is not unique... you could've also used the other (9, -18).... :)

OpenStudy (anonymous):

*other point (9, -18)

OpenStudy (anonymous):

thank you so much! do you mind helping?

OpenStudy (anonymous):

i have one more

OpenStudy (anonymous):

go ahead and post up a new question.... i don't think open study will run out of hard drive space anytime soon.

OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5

OpenStudy (anonymous):

this is standard form: Ax + By = C, where A, B, C are integer coefficients... so in your slope-intercept form, we need to clear the decimals by multiplying by the whole equation by 10... \(\Large y = –1.7x + 8.5 \) \(\Large \color {red}{10}[y = –1.7x + 8.5] \) can you tell me what youd get here?

OpenStudy (anonymous):

17x + 10y = 85 ?

OpenStudy (anonymous):

u got it.... :) nice job..